OpenStudy (anonymous):

Tanya and Akira wrote the point-slope form of an equation for a line that passes through (-2,-6) and (1,6). Tanya says that Akira's equation is wrong . Akira says they are both correct. Help please!!

OpenStudy (anonymous):

@kidrah69

OpenStudy (kidrah69):

What were the equations?

OpenStudy (anonymous):

Oh opps sorry! Tanya: y+6=4(x+2) & Akira: y-6=4(x-1)

OpenStudy (kidrah69):

Ok we need to find the slope first. \[\frac{ y2-y1 }{ x2-x1 }=slope \] plug in the numbers and solve.

OpenStudy (anonymous):

Oh wait I think I've got it now. Thanks so much!

OpenStudy (kidrah69):

:)